OpenStudy (anonymous):

(x+4)/6 = 5/6 - (x-1)/4 Please help. I worked this out already but got it wrong.

OpenStudy (tennis5518):

hope that help @katiexo

OpenStudy (anonymous):

still lost. thought I had to find the LCD first and get a common denominator

OpenStudy (tennis5518):

yea nvm i make it easy

OpenStudy (anonymous):

i know the lcd is 12

OpenStudy (tennis5518):

x+46=56−x−14 ok so Step-By-Step Solution: Multiply both sides of the equation by 6. (x+4)=(56−x−14)⋅(6)

OpenStudy (mathmale):

Indeed, finding the lowest common denominator and then using it to eliminate the fractions is the way to go. I agree that 12 is the LCD. Good. What next? Why not copy the problem, as given, and then mult. each term by the LCD? Simplify the result. Good-bye, fractions.

OpenStudy (anonymous):

I get to 10 - 3x - 3 = 2x + 8, but when I simplify further I get the wrong answer.

OpenStudy (anonymous):

@mathmale

RhondaSommer (rhondasommer):

(x+4)/6 = 5/6 - (x-1)/4 (x+4)/6 = 5/6 + (-x+1)/4 (x+4) = 5 + (- 6x + 6)/4 (multiplied both sides by 6) 4x + 16 = 5 - 6x +6 (multiplied both sides by 4) 4x + 16 = -6x +11 (simplified 5+6) can you simplify that?

OpenStudy (mathmale):

Indeed, finding the lowest common denominator and then using it to eliminate the fractions is the way to go. I agree that 12 is the LCD. Good. What next? Why not copy the problem, as given, and then mult. each term by the LCD? Simplify the result. Good-bye, fractions. Could we please come to agreement? Is the LCD 6 or is it 12?

OpenStudy (anonymous):

It's 12, that's why I multiplied the terms by 12

OpenStudy (mathmale):

Great. And your results?

OpenStudy (anonymous):

Then I worked it out and got \[\frac{ 10 }{ 12 } - \frac{ 3x-3 }{ 12 } -\frac{ 2x+8 }{ 12 } = 0\]

RhondaSommer (rhondasommer):

then I double checked my answer

OpenStudy (anonymous):

My answer key has the correct answer as simply 1

RhondaSommer (rhondasommer):

but it isnt...

OpenStudy (anonymous):

This is the original equation (Ignore the '3')

OpenStudy (mathmale):

Ah, a little friendly competition! Katie: Now that you have the same denom.., cancel out that den. What's left of your equation? Please show all the steps you go through to find the value of x that satisfies this equation.

RhondaSommer (rhondasommer):

how would 1 work?

OpenStudy (anonymous):

So after that I just set the numerators equal to each other because they have the same denominator. So I was left with \[10-3x-3-2x+8 = 0\]

OpenStudy (anonymous):

@RhondaSommer where are you getting that equation from? that's not the original equation. check my picture attachment

OpenStudy (mathmale):

And? Combine like terms and solve for x.

OpenStudy (mathmale):

Remember, you'll have to check your own answer by substitution.

RhondaSommer (rhondasommer):

lol I simply got myself confused...mathmale is right XD

OpenStudy (anonymous):

Then I got \[15-5x = 0\]

RhondaSommer (rhondasommer):

so sorry XD

OpenStudy (anonymous):

Then ended with x = 3 but that's not the answer in my answer key :/

OpenStudy (anonymous):

I'm just so confused!

OpenStudy (anonymous):

I plugged it back in and it's not right either. No idea where I'm going wrong...

OpenStudy (mathmale):

Relax. You've demonstrated that you know what "Lcd" means and that you know how to find it. You do have an arithmetic error that we can fix.

OpenStudy (anonymous):

I can't find my error

OpenStudy (mathmale):

Your original problem: screenshot20160207at6.08.30pm.png

OpenStudy (mathmale):

Could you post it yoursxelf, katie?

OpenStudy (mathmale):

You've stated, correctly, that the LCD is 12. Mult. the first fraction of your original eq'n by 12 and simplify the result.

OpenStudy (anonymous):

Yes. It's \[\frac{ 5 }{ 6 } - \frac{ x-1 }{ 4 } = \frac{ x+4 }{ 6 }\]

OpenStudy (mathmale):

Mult the first term by 12 and simplify the result as much as possible. What do you get?

OpenStudy (mathmale):

Hint: 12/6 = 2, and 2*5+?

OpenStudy (mathmale):

xxxx and 2*5=??

OpenStudy (anonymous):

10/12 which reduces to 5/6?

OpenStudy (mathmale):

Katie, we found the LCD and need to use it to eliminate fractions. So, no, we don't want that 5/6. 2*5=?

OpenStudy (anonymous):

10

OpenStudy (mathmale):

Yes. Now mult the 2nd term by 12 and reduce the result. Result?

OpenStudy (anonymous):

\[\frac{ 3x-3 }{ 12 } ?\]

OpenStudy (mathmale):

Why the 12 in the denom.? Again, our purpose was to eliminate fractions. That has not happened yet.

OpenStudy (anonymous):

Because I was trying to get all the denoms the same first

OpenStudy (anonymous):

So just 3x-3?

OpenStudy (mathmale):

The 2nd term is -(x-1)/4. Mult that by 12 and cancel the 4. You'd be correct if you'd please enclose that 3x-3 in parentheses and put a neg. sign in front of it. Try again.

OpenStudy (mathmale):

The 2nd term is -(x-1)/4. Mult that by 12 and cancel the 4.

OpenStudy (anonymous):

-(3x-3)?

OpenStudy (mathmale):

Remember: the purpose of the LCD is to enable us to combine fractions.

OpenStudy (mathmale):

In this particular case we also wanted to eliminate fractions. Mult the 3rd term by 12 and simplify the result, please.

OpenStudy (anonymous):

(2x+8)

OpenStudy (mathmale):

Your -(3x-3) is fine, by the way. Good. Now reconstruct the equation with NO denominator.

OpenStudy (mathmale):

10 - ?? = ??

OpenStudy (anonymous):

10 - (3x-3) = (2x+8)

OpenStudy (mathmale):

Great. Now, carefully, eliminate the parentheses.

OpenStudy (anonymous):

10 - 3x+3 = 2x+8

OpenStudy (mathmale):

Combine like terms. The 2x on the right is already positive, so leave it there and move the 3x term.

OpenStudy (anonymous):

13 = 5x + 8

OpenStudy (mathmale):

subtract 8 from both sides and siimplify the equation.

OpenStudy (anonymous):

x = 1!!!!

OpenStudy (anonymous):

Oh my gosh finally!!! I see my only problem was that -(3x-3).

OpenStudy (mathmale):

Now, go back to the original equation. Subst. 1 for x. Is the equation now true?

OpenStudy (anonymous):

Yes! It'd be 5/6 = 5/6 - 0, so they're equal

OpenStudy (mathmale):

Therefore, your x=1 is the correct solution. Congrats.

OpenStudy (anonymous):

Thanks for all your help! :)

OpenStudy (mathmale):

Mygreat pleasure, Katie. See you!